s, Interest, inflation, and purchasing power Suppoce Ana is an avid reader and buys only comic books. Ana depasits $4,000 In a bank account that pays an anmual nominal interest rate of 15%. Assume this interest rate is fixed-that is, it wont change over time. At the time of her deposit, a comic book is priced at $20.00. nat 4,000 deposit is). 200 comic books. For sach of the annuel infistion rates glven in the fowowing inflation. Then enter the value for the real interest rate at each of the given ind 100 ces. o table, first determine the new price of a comie book, assuming kt reses at the rate of enter the corresponding purchasing power of Angs dengsit after ane vear in the first row of tho table for aach Jotiation rate. e first row down to the nearest 200 ook. For example, # you find that the deposit will over 20.7 comic books, you would round the p rchasing power down to 20 comic books und ssumption that Ana will not buy seven-tenths of s comic book. 230 Annual Inflation Rate 388 15% Number of Comics Ana Can Purchase after One Year Real Interest Rate When the rate of inflation is greater than the interest rate ce anas deposit, the purchasing pewer of her deposit course of the year tells w over the
Answer #1

Initially, the purchasing power of Ana's $ 4000 deposit is = 4000/20 = 200 comic books

Value of deposit after one year = 4000 x (1 + 15%) = $ 4600

Annual inflation rate
0% 15% 18%
New price 20 23 23.6
Number of comics 230 200 194

New price = 20 x (1 + inflation rate)

Number of comics = 4600/New price

When the rate of inflation is greater than the interest rate on Ana's deposit, the purchasing power of her deposit falls over the course of the year.